# Demystifying 5G Polar and LDPC Codes: A Comprehensive Review and Foundations

### v2

@@ Abstract
This paper serves as a comprehensive guide for practitioners and scholars aiming to understand the channel coding and decoding schemes integral to the 5G NR standard, with a particular focus on LDPC and polar codes. We start by explaining the design procedures that underlie these channel codes, offering fundamental information from perspectives of both encoding and decoding. In order to determine the present status of research in this area, we also provide a thorough literature review. Notably, we add comprehensive, standard-specific information to these foundational evaluations that is frequently difficult to extract from technical specification documents. The significance of reviewing and refining the foundations of the aforementioned codes lies in their potential to serve as candidate error-correcting codes for the future 6G standard and beyond.
@@ I Introduction
The Channel coding emerged when Shannon applied probability theory to the study of communication systems Shannon [1]. In his seminal work , Shannon demonstrated that errors introduced by a noisy channel can be reduced to a desired level through the use of an appropriate coding scheme, for a transmission rate that is less than or equal to the channel’s capacity. To optimize the performance of communication channels, various techniques can be employed to mitigate the effects of noise, disturbances, or distortions, while increasing the channel throughput until it approaches its maximum capacity. This can be achieved either by physically modifying the signal through modulation or by altering its content through channel coding. Most error-correcting codes are linear codes, which, due to their systematic representation, are classified as systematic codes. Linear block codes can be understood in a variety of ways, and some of the most commonly used in many contexts are Reed Solomon (RS) codes [2], Hamming codes [3], CRC codes, Bose–Chaudhuri–Hocquenghem (BCH) codes [4] and LDPC codes [5]. It is possible to represent a linear block code in matrix form or in the form of a graph factor.
  • Generator Matrix: $\mathbf{G}$ is the generator matrix of dimension $\mathsf{K}\times\mathsf{N}$. The matrix is said to be systematic if $\mathbf{G}=[\mathbf{P}\ \mathbf{I}_{\mathsf{K}}]$ where $\mathbf{I}_{\mathsf{K}}$ and $\mathbf{P}$ is called the parity matrix of dimension ($\mathsf{K}\times\mathsf{N-K}$). In this case, the code is formed by parity bits followed by message bits.
    The matrix $\mathbf{G}$ is said to be non-systematic if $\mathbf{G}=[\mathbf{P}\ \mathbf{M}]$ where $\mathbf{M}$ is any matrix of size ($\mathsf{K}\times\mathsf{K}$).
  • Parity Check Matrix: $\mathbf{H}$ is called the parity check matrix denoted as $\mathbf{H}=[\mathbf{I}_{\mathsf{N-K}}\ \mathbf{P}^{\mathsf{T}}]$, where $\mathbf{P}^{\mathsf{T}}$ is the transpose of the parity matrix $\mathbf{P}$. Thus, it can be easily shown that $\mathbf{G}\cdot\mathbf{H}^{\mathsf{T}}=\mathbf{0}$, where operations are performed in $\mathbb{F}_{2}$.
  • Syndrome Checking: For each code word $\mathbf{c}$: $\mathbf{c}\cdot\mathbf{H}^{\mathsf{T}}=\mathbf{m}\cdot\mathbf{G}\cdot\mathbf{H}^{\mathsf{T}}=\mathbf{0}$, where $\mathbf{m}=[\mathrm{m}(0),\mathrm{m}(1),\ldots,\mathrm{m}(\mathsf{K}-1)]$. If $\mathbf{c}\cdot\mathbf{H}^{\mathsf{T}}\neq\mathbf{0}$, it means there are errors in the received bits.
In 5G NR, data and control streams from PHY to MAC layer are encoded /decoded to offer transport and control services over the radio transmission link. Channel coding scheme is a combination of error detection, error correcting, rate matching, interleaving and transport channel or control information mapping onto/splitting from physical channels.
Table I and Table II show the usage of coding scheme for the different types of Transport channels and for the different control information types [6].
@@
TABLE I: Usage of channel coding scheme for transport channels (TrCHs). [6]
@@
UL-SCH LDPC
DL-SCH LDPC
PCH Polar code
BCH Polar code
@@
The article is structured as follows. Section II and Section III lays out the foundations and detailed reviews of 5G NR LDPC codes and 5G NR Polar codes respectively, Section IV presents the numerical results and performance analysis, and finally Section V concludes the paper.
@@
PUSCH
Physical Uplink Shared Channel
QPSK
Quadrature Phase Shift Keying
@@ II-B Foundations and Fundamentals
The LDPC code is presented in matrix form $[\mathbf{G},\mathbf{H}]$, where $\mathbf{G}$ is the generator matrix of the code and $\mathbf{H}$ is the parity check matrix (PCM). The parity matrix $\mathbf{H}$ is sparse, containing very few ones. It can be represented in the form of a Tanner graph. This graph consists of two types of nodes: bit nodes (BNs) and check nodes (CNs) which are connected by edges. The BNs and CNs correspond, respectively, to the columns (code word bits) and rows (parity constraints) of the matrix $\mathbf{H}$. A variable node $i$ is connected to a check node $j$ if $\mathbf{H}(i;j)=1$. This graph is bipartite, since nodes of the same type cannot be connected (i.e., a CN cannot be connected to another CN). Tanner graphs are commonly used to represent the parity matrix in LDPC codes. A Tanner graph leads to decoding algorithms of fairly low complexity [7].
@@
NR LDPC code is a family of QC-LDPC codes. It is constructed from a matrix named $\mathbf{H}_{\mathsf{BG}}$ of dimension $M\times N$ called base graph matrix $\mathbf{BG}$. The $\mathbf{H}_{\mathsf{BG}}$ matrices are selected in the 5G NR coding process according to the coding rate and the length of the transport block or code block. Thus, for $\mathbf{BG}_{\mathsf{1}}$ ($N=68$, $M=46$) and for $\mathbf{BG}_{\mathsf{2}}$ ($N=52$, $M=42$). Since $\mathbf{BG}_{\mathsf{1}}$ is targeted for larger block length $\mathsf{K}\leq 8448$ and coding rates between $91/3\leq\mathrm{R}\leq 8/$ 9, $\mathbf{BG}_{\mathsf{2}}$ is employed for small blocks $\mathsf{K}\leq 3840$ and coding rates between $31/5\leq\mathrm{R}\leq 2/$ 3.
@@
For $\mathbf{BG}_{\mathsf{1}}$, $\mathbf{A}$ is a $4\times 10$ matrix; $\mathbf{E}$ is a $4\times 4$ matrix; 00 is $4\times 38$ all zero matrix; $\mathbf{B}$ is a $38\times 10$ matrix; $\mathbf{C}$ is a $38\times 4$ matrix; $\mathbf{I}$ is $38\times 38$ identity matrix; Sub-matrix $\mathbf{E}$ is a double diagonal matrix that is benefit for encoding. An example of $\mathbf{BG}_{\mathsf{1}}$ with set index of listing size (iLS) =$1$ in 3GPP TS 38.212[6] standard is shown in Figure 2. In order to distinguish with the number $1$ in base graphs in 3GPP TS 38.212 [6] standard, null value in the base graph will be replaced by $-1$.
@@
Set index (iLS) Set of lifting sizes (Z)
0 2, 4, 8, 16, 32, 64, 128, 256
1 3, 6, 12, 24, 48, 96, 192, 384
2 5, 10, 20, 40, 80, 160, 320
3 7, 14, 28, 56, 112, 224
4 9, 18, 36, 72, 144, 288
5 11, 22, 44, 88, 176, 352
6 13, 26, 52, 104, 208
7 15, 30, 60, 120, 240
@@
  • Each element of value $-1$ in $\mathbf{H}_{\mathsf{BG}}$ is replaced by a null matrix of size $\mathrm{Z_{c}}\times\mathrm{Z_{c}}$.
  • Each element of value 00 in $\mathbf{H}_{\mathsf{BG}}$ is replaced by an identity matrix $\mathbf{I}$ of size $\mathrm{Z_{c}}\times\mathrm{Z_{c}}$.
@@
$\mathbf{H}_{\mathsf{BG}}=\left[\begin{array}[]{cccccc}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}2}&{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-1}&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}1}&{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}3}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}&{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-1}\\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}1}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}&{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-1}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-1}&{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}3}&{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}2}&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}1}&{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-1}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}0}\end{array}\right],$ (2)
@@ 2 $B$leq192 6
    The rate matching aims to adapt different code rates. Rate matching is based on redundancy version (RV) from 0 to 3 [6]. Each RV divides the base graph, excluding the first two columns, into four chunks at different positions. Note that the first two columns are always punctured to improve performance. RV 00 is well suited for the first transmission and has good self-decodability [44].
    Hence, the rate matching is carried out on each code block independently. Assume that the coded message bit output from the LDPC encoder of the $r-th$ code block is $d_{r}(1),d_{r}(2),\ldots,d_{r}(\mathrm{N}_{r})$ and $E_{r}$ is the length of the output message after performing rate matching on this $r-th$ coded message. Thus, the output bit-message from the rate matcher is denoted by $e_{r}(1),e_{r}(2),\ldots,e_{r}(E_{r})$ which is calculated using the following equation:
@@
    The CB de-concatenation is used to break the transport block message into C numbers of code blocks message. Assume that the input message to code block de-concatenation is $y(0),y(1),\ldots,y(G-1)$. The output message from code block de-concatenation is $f_{r}(0),f_{r}(1),\ldots,f_{r}(E_{r}-1)$
@@ II-D Soft-Decision based LDPC decoding Algorithms
Figure 4: Check Node Updates: $\mathsf{CN}_{4}\longrightarrow\mathsf{BN}_{5}$
@@
Index sets $\mathsf{BN}_{j\backslash i}$ and $\mathsf{CN}_{i\backslash j}$ are based on the parity check matrix (PCM). Index set $\mathsf{CN_{i}}$ and $\mathsf{BN_{j}}$ correspond to all non-zero element on column $i$ and row $j$ of the PCM, respectively. Figure 6 a simple conceptual illustration of BN and CN index sets within the PCM provided in (1) for the specified values of $i=3$ and $j=2$.
@@
  • $\mathsf{BN}_{j\backslash i}$ =bit nodes connected to check node $j$, excluding bit node $i$,
@@
  • $\mathsf{CN}_{i\backslash j}$ =check nodes connected to bit node $i$, excluding check node $j$,
@@
  • The check-to-variable extrinsic message passing to the $j-th$ variable node from the $i-th$ check node is denoted by $\mathscr{L}\left(\mathrm{r}_{ji}\right)$ and $\mathscr{L}\left(\mathrm{q}_{ij}\right)$ is the variable-to-check extrinsic message.
@@
Where $\phi(\mathbf{x})=-\log\left[\tanh\left(\frac{|\mathbf{x}|}{2}\right)\right]=\log\left(\frac{\mathrm{e}^{\mathbf{x}}+1}{\mathrm{e}^{\mathbf{x}}-1}\right)$
@@
$\mathscr{L}\left(\mathrm{q}_{ij}\right)=\mathscr{L}\left(\mathrm{c}_{i}\right)+\sum_{j^{\prime}\in\mathsf{CN}_{i\backslash j}}\mathscr{L}\left(\mathrm{r}_{j^{\prime}i}\right).$ (20)
@@
$\hat{\mathrm{c}}_{i}=\left\{\begin{array}[]{lc}1,&\text{ if }\mathscr{L}\left(Q_{i}\right)<0,\\ 0,&\text{ else .}\end{array}\right.$ (22)

### v3

@@ Abstract
Understanding how 5G networks correct errors is no trivial matter. At the heart of the process lie two sophisticated families of codes: LDPC and polar codes. This paper opens the black box, not only by explaining how these codes are designed, but also by showing how they are encoded and decoded in practice. To map where research currently stands, we present a detailed survey of the literature supplemented with insights that are often buried deep within technical standards. These foundations are not just historical footnotes: they are strong candidates for powering error correction in 6G and beyond. In bringing clarity to these building blocks, we aim to help engineers and researchers navigate what is both a complex and increasingly vital part of wireless communication.
@@ I Introduction
C hannel coding was born the day Shannon showed that even noisy communication could be made nearly error-free—provided we’re willing to encode wisely. His 1948 paper laid the foundation: by embedding redundancy into transmitted data, one could correct errors as long as the data rate stayed below a theoretical threshold, now known as the channel capacity. Since then, the quest has been clear: how do we push transmission rates as close as possible to that limit while still taming noise, distortion, and interference? Two tools remain essential: modulation, which shapes the physical signal, and channel coding, which reshapes the message itself. Most modern codes are linear and systematic, meaning their structure directly reflects the message they carry. Among the many varieties, Reed–Solomon[2], Hamming[3], BCH [4], CRC, and LDPC [5] codes have emerged as cornerstones—each bringing unique properties suited to different applications. These codes can be expressed algebraically, via matrices, or visually, as factor graphs.
It is possible to describe a linear block code either algebraically, using matrices, or structurally, through graphical models such as factor graphs. In practice, the matrix-based representation is most common, relying on two fundamental constructs: the generator matrix and the parity-check matrix.
A linear block code $\mathcal{C}$ of length $\mathsf{N}$ and dimension $\mathsf{K}$ over the binary field $\mathbb{F}_{2}$ is defined as the image of a generator matrix $\mathbf{G}\in\mathbb{F}_{2}^{\mathsf{K}\times\mathsf{N}}$:
$\mathcal{C}=\left\{\mathbf{c}=\bm{m}\cdot\mathbf{G}\mid\bm{m}\in\mathbb{F}_{2}^{\mathsf{K}}\right\}.$ (1)
A generator matrix $\mathbf{G}$ is said to be in systematic form if it can be expressed as: $\mathbf{G}=\left[\mathbf{I}_{\mathsf{K}}\;\;\mathbf{P}\right]$, where $\mathbf{I}_{\mathsf{K}}$ denotes the $\mathsf{K}\times\mathsf{K}$ identity matrix and $\mathbf{P}\in\mathbb{F}_{2}^{\mathsf{K}\times(\mathsf{N}-\mathsf{K})}$ is the associated parity matrix.
More generally, $\mathbf{G}$ need not expose the identity structure and may be expressed in non-systematic form as: $\mathbf{G}=\left[\mathbf{P}\;\;\mathbf{M}\right]$, where $\mathbf{M}\in\mathbb{F}_{2}^{\mathsf{K}\times\mathsf{K}}$ is any invertible matrix.
The code $\mathcal{C}$ can alternatively be specified by a parity-check matrix $\mathbf{H}\in\mathbb{F}_{2}^{(\mathsf{N}-\mathsf{K})\times\mathsf{N}}$, satisfying: $\mathcal{C}=\left\{\mathbf{c}\in\mathbb{F}_{2}^{\mathsf{N}}\;\middle|\;\mathbf{H}\cdot\mathbf{c}^{\mathsf{T}}=\mathbf{0}\right\}$.
[…]
Given a received vector $\bm{r}\in\mathbb{F}_{2}^{\mathsf{N}}$, its syndrome is computed as: $\bm{s}=\mathbf{H}\cdot\bm{r}^{\mathsf{T}}$. If $\bm{s}=\mathbf{0}$, then $\bm{r}$ is a valid codeword (assuming no transmission error). Otherwise, $\bm{s}\neq\mathbf{0}$ indicates the presence of errors, and the syndrome can be used to infer the error pattern.
@@
TABLE I: Usage of channel coding scheme for transport channels (TrCHs)[6].
@@
LDPC UL-SCH
DL-SCH
Polar code PCH
BCH
In 5G New Radio (NR), these principles take on practical form. Information whether data or control is encoded as it flows from the physical layer (PHY) to the MAC, enabling robust communication over unpredictable wireless channels. The coding scheme used in 5G is a carefully engineered blend of error detection, correction, rate matching, interleaving, and mapping onto physical resources. Tables I and II summarize the coding choices made in the 5G NR standard for various transport and control channels [6].
@@
The article is structured as follows. Section II and Section III lays out the foundations and detailed reviews of 5G NR LDPC codes and 5G NR Polar codes respectively, and finally Section IV concludes the paper.
@@ II-B Foundations and Fundamentals
The LDPC code is presented in matrix form $\mathbf{G},\mathbf{H}]$, where $\mathbf{G}$ is the generator matrix of the code and $\mathbf{H}$ is the parity check matrix (PCM). The parity matrix $\mathbf{H}$ is sparse, containing very few ones. It can be represented in the form of a Tanner graph. This graph consists of two types of nodes: bit nodes (BNs) and check nodes (CNs) which are connected by edges. The BNs and CNs correspond, respectively, to the columns (code word bits) and rows (parity constraints) of the matrix $\mathbf{H}$. A variable node $i$ is connected to a check node $j$ if $\mathbf{H}(i;j)=1$. This graph is bipartite, since nodes of the same type cannot be connected (i.e., a CN cannot be connected to another CN). Tanner graphs are commonly used to represent the parity matrix in LDPC codes. A Tanner graph leads to decoding algorithms of fairly low complexity [7].
@@
NR LDPC code is a family of QC-LDPC codes. It is constructed from a matrix named $\mathbf{H}_{\mathsf{BG}}$ of dimension $M\times N$ called base graph matrix $\mathbf{BG}$. The $\mathbf{H}_{\mathsf{BG}}$ matrices are selected in the 5G NR coding process according to the coding rate and the length of the transport block or code block. Thus, for $\mathbf{BG}_{\mathsf{1}}$ ($N=68$, $M=46$) and for $\mathbf{BG}_{\mathsf{2}}$ ($N=52$, $M=42$). Since $\mathbf{BG}_{\mathsf{1}}$ is targeted for larger block length $\mathsf{K}\leq 8448$ and coding rates between $1/3\leq\mathrm{R}\leq 8/9$, $\mathbf{BG}_{\mathsf{2}}$ is employed for small blocks $\mathsf{K}\leq 3840$ and coding rates between $1/5\leq\mathrm{R}\leq 2/3$.
@@
For $\mathbf{BG}_{\mathsf{1}}$, $\mathbf{A}$ is a $4\times 10$ matrix; $\mathbf{E}$ is a $4\times 4$ matrix; 0 is $4\times 38$ all zero matrix; $\mathbf{B}$ is a $38\times 10$ matrix; $\mathbf{C}$ is a $38\times 4$ matrix; $\mathbf{I}$ is $38\times 38$ identity matrix; Sub-matrix $\mathbf{E}$ is a double diagonal matrix that is benefit for encoding. An example of $\mathbf{BG}_{\mathsf{1}}$ with set index of listing size (iLS) =$1$ in 3GPP TS 38.212[6] standard is shown in Figure 2. In order to distinguish with the number $1$ in base graphs in 3GPP TS 38.212 [6] standard, null value in the base graph will be replaced by 1-1.
@@
. Set index (iLS) Set of lifting sizes (Z) 0 2, 4, 8, 16, 32, 64, 128, 256 1 3, 6, 12, 24, 48, 96, 192, 384 2 5, 10, 20, 40, 80, 160, 320 3 7, 14, 28, 56, 112, 224 4 9, 18, 36, 72, 144, 288 5 11, 22, 44, 88, 176, 352 6 13, 26, 52, 104, 208 7 15, 30, 60, 120, 240
@@
  • Each element of value 1-1 in $\mathbf{H}_{\mathsf{BG}}$ is replaced by a null matrix of size $\mathrm{Z_{c}}\times\mathrm{Z_{c}}$.
  • Each element of value 0 in $\mathbf{H}_{\mathsf{BG}}$ is replaced by an identity matrix $\mathbf{I}$ of size $\mathrm{Z_{c}}\times\mathrm{Z_{c}}$.
@@
[…]
@@ 2 $B$ $\leq$192 6
    The rate matching aims to adapt different code rates. Rate matching is based on redundancy version (RV) from 0 to 3 [6]. Each RV divides the base graph, excluding the first two columns, into four chunks at different positions. Note that the first two columns are always punctured to improve performance. RV 0 is well suited for the first transmission and has good self-decodability [44].
    Hence, the rate matching is carried out on each code block independently. Assume that the coded message bit output from the LDPC encoder of the rthr-th code block is $d_{r}(1),d_{r}(2),\ldots,d_{r}(\mathrm{N}_{r})$ and $E_{r}$ is the length of the output message after performing rate matching on this rthr-th coded message. Thus, the output bit-message from the rate matcher is denoted by $e_{r}(1),e_{r}(2),\ldots,e_{r}(E_{r})$ which is calculated using the following equation:
@@
    The CB de-concatenation is used to break the transport block message into C numbers of code blocks message. Assume that the input message to code block de-concatenation is $y(0),y(1),\ldots,y(G-1)$. The output message from code block de-concatenation is $f_{r}(0),f_{r}(1),\ldots,f_{r}(E_{r}-1)$ .
@@ II-D Soft-Decision based LDPC decoding Algorithms
Figure 4: Check Node Updates: $\mathsf{CN}_{4}\longrightarrow\mathsf{BN}_{5}$ .
@@
[…]
@@
[…]
@@
[…]
@@
  • The check-to-variable extrinsic message passing to the jthj-th variable node from the ithi-th check node is denoted by $\mathscr{L}\left(\mathrm{r}_{ji}\right)$ and $\mathscr{L}\left(\mathrm{q}_{ij}\right)$ is the variable-to-check extrinsic message.
@@
Where $\phi(\mathbf{x})=-\log\left[\tanh\left(\frac{|\mathbf{x}|}{2}\right)\right]=\log\left(\frac{\mathrm{e}^{\mathbf{x}}+1}{\mathrm{e}^{\mathbf{x}}-1}\right)$ .
@@
[…]
@@
[…]
@@
1 Input: The channel log likelihoods : $\mathscr{L}_{i}\in\mathbb{R}^{n}$:
2 Maximum # of iterations, MAXITER
3 Description of the parity check matrix using $\mathsf{BN}(j)$ and $\mathsf{CN}(i)$.
[…]
5 Initialization :
6
7for each i, and for each j $\in\mathsf{CN}(i)$  do
8 $\mathscr{L}\left(\mathrm{q}_{ij}\right)=\mathscr{L}_{i}$
9Check Node to Variable Node Step (horizontal step):
10 for each check node $j$ do
11 for each variable node $i$ $\in$ $\mathsf{BN}(j)$ do
[…]
13
14Variable Node to Check Node Step (vertical step)
15 for each variable node $i$ do
16 for each check node $j$ $\in$ $\mathsf{CN}(i)$ do
[…]
18    Also compute the output likelihoods
19    $\mathscr{L}\left(\mathrm{Q}_{i}\right)=\mathscr{L}_{i}+\displaystyle\sum_{j^{\prime}\in\mathsf{CN}_{i}}\mathscr{L}\left(\mathrm{r}_{j^{\prime}i}\right)$
20
21Hard decision:
22 for each $i$ do
23 $\hat{\mathrm{c}}_{i}=1$ if $\mathscr{L}\left(\mathrm{Q}_{i}\right)<0$ else $\hat{\mathrm{c}}_{i}=0$
24Parity Check:
25 if  $\mathbf{H\hat{c}^{\mathsf{T}}}=\mathbf{0}$ then
26  return $\mathbf{\hat{c}}$
27otherwise, $\mathbf{if}$ # ITER << MAXITER
28 goto Check Node to Variable Node Step
29 else
30  return $\mathbf{\hat{c}}$ and indication of coding failure.
Algorithm 1 Log-Likelihood Belief Propagation Decoding Algorithm